if 9x=4 what does 9-2x equals

if 9x=4 what does 9-2x equals
Answer #1

oma me if Gc=4 what_does 9-222 eguals Solution: - 9XE4 . X=4 1 put 속 . - 2x = 9-12) 9-3 1 1-8 ។ 73 9-22 = 13 = 8.1)